OpenStudy (anonymous):

How would I go about taking the double derivative of the function F(x) = integral[1, 3x] ln(t^2)dt? The only problem I have is with the initial derivative.

hartnn (hartnn):

\(\Large \dfrac{d}{dx}\int \limits_{a}^{g(x)}h(x)dx = h(g(x))g'(x) \)

OpenStudy (anonymous):

So F'(x) would be 3ln((3x)^2)?

hartnn (hartnn):

yes!

hartnn (hartnn):

you can simplify that further

OpenStudy (anonymous):

3ln(9(x^2))?

hartnn (hartnn):

\(\ln A^2 = 2 \ln A\)

OpenStudy (anonymous):

6ln(3x)? or would it be 6ln(9x)?

OpenStudy (anonymous):

Oh, 6ln(3x).

hartnn (hartnn):

yes

OpenStudy (anonymous):

Thank you so much for the help.

hartnn (hartnn):

need help in finding its derivative? whats the final answer you got? just for verification

OpenStudy (anonymous):

6/x

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

Thanks.

hartnn (hartnn):

welcome ^_^